OpenStudy (anonymous):

Can someone help on this trig proof? (1+cosx)/(1-cosx)=(cscx+cotx)^2 I started by eliminating cos from the denominator Multiply by ((1+cosx)/(1+cosx)) (1+2cosx+cos^2x)/(1-cos^2x) use identity to change 1-cos^2x to sin^2x so now I have (1+2cosx+cos^2x)/sin^2x I am stuck. Did I do a step wrong?

OpenStudy (solomonzelman):

Are you allowed to play with both sides at the same time, or not?

OpenStudy (anonymous):

no I have to work from one side to equal the other

OpenStudy (anonymous):

if you can reverse engineer it I think that is vaild :)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{ 1+\cos x }{1-\cos x} =\left(\csc x + \cot x\right)^2 }\) \(\large\color{black}{ \displaystyle \frac{ 1+\cos x }{1-\cos x} =\left(\frac{1}{\sin x} + \frac{\cos x}{\sin x}\right)^2 }\) \(\large\color{black}{ \displaystyle \frac{ 1+\cos x }{1-\cos x} =\left(\frac{1+\cos x}{\sin x}\right)^2 }\) \(\large\color{black}{ \displaystyle \frac{ 1+\cos x }{1-\cos x} =\frac{(1+\cos x)^2}{\sin^2x} }\) \(\large\color{black}{ \displaystyle \frac{ 1+\cos x }{1-\cos x} =\frac{(1+\cos x)^2}{1-\cos^2x} }\)

OpenStudy (solomonzelman):

then, the last step (on the right side), is to divide by \(1+\cos x\) on top and bottom.

OpenStudy (anonymous):

Oh that is so much more simplistic! Beautiful Solomon!

OpenStudy (solomonzelman):

Anytime:)